# SPEAR: A Unified SSL Framework for Learning Speech and Audio Representations

### v1

@@ 4 Experimental Setup
Data
Pre-training is performed on a mixture of public unlabelled English speech datasets and general audio datasets, as shown in Table 1. Due to the limited amount of public general audio datasets, we incorporate two music datasets, Music4all (Santana et al., 2020) and MTG-Jamendo (Bogdanov et al., 2019), to enrich the general audio data.
Model Architecture
As shown in (Shi et al., 2024), modelling speech representations at different time resolutions is beneficial to the comprehensive capabilities of speech SSL models. Therefore, Zipformer (Yao et al., 2024) is selected as the feature encoder in SPEAR due to its dynamic downsampling mechanism in the intermediate layers. The model receives 128-dimensional filter-bank features as input and produces frame-level representations at a 50 Hz frame rate.
| Model | Domain(s) | Data Mixture | Total Hours |
@@
| SPEAR-{Base, Large} | Speech & Audio | Mix-97k | 97k |
| SPEAR XLarge | Speech & Audio | Mix-197k | 197k |
@@
We pre-train three model scales under the SPEAR framework: Base (94M), Large (327M), and XLarge (600M). At the Base and Large scales, we pre-train both single-domain and dual-domain models. To further explore the benefits of scaling, we additionally train an XLarge dual-domain model on a larger dataset. The pre-training data used for different scales of SPEAR models are given in Table 2. The data mixtures are defined as follows: Speech-84k comprises Libriheavy, GigaSpeech, and VoxPopuli (en); Audio-13k includes all general audio datasets listed in Table 1; Mix-97k combines Speech-84k and Audio-13k; and Mix-197k additionally includes the Yodas-granary dataset. Detail regarding the encoder configurations is provided in Appendix C.1, and hyperparameters for pre-training are presented in Appendix C.2.
| Teacher Model | # Params |
Model Config | MVQ Config | |||||
| Domain | Model Dim | Frame Rate | |||||||
| WavLM Large | 317M | 94k | Speech | 1024 | 50 Hz | 16 | 256 | ||
| Dasheng 1.2B | 1.2B | 272k | Audio | 1536 | 25 Hz | 8 | 256 | ||
Teacher Models and MVQ quantiser
WavLM Large (Chen et al., 2022) and Dasheng 1.2B (Dinkel et al., 2024) are utilised to generate pre-training targets for speech and audio domains, respectively. Model details and corresponding MVQ quantiser configurations are provided in Table 3. WavLM Large is pre-trained on 94k hours of unlabelled English speech data, including Libri-light (Kahn et al., 2020), GigaSpeech, and Voxpopuli (en). It can be fairly contrasted with SPEAR models trained on Speech-84k since Libriheavy is the segmented version of Libri-light. The speech MVQ quantiser with 16 codebooks is trained using the 21st layer representations of WavLM Large on 100 hours of data sampled from LibriSpeech (Panayotov et al., 2015). Dasheng 1.2B is an audio SSL model pre-trained with an MAE (Huang et al., 2022) objective on an exceptionally large amount of general audio data of over 272k hours. The audio MVQ quantiser is trained on the last-layer representations with 8 codebooks on 50 hours of AudioSet balanced set.
5 Results
To validate the effectiveness of the SPEAR framework, we assess its performance through both full fine-tuning and frozen representation evaluations on two major benchmarks for evaluating speech and audio representations: SUPERB Yang et al. (2021); Tsai et al. (2022) and HEAR (Turian et al., 2022). Additionally, extensive ablation studies on the core components of SPEAR are performed in Appendix G to provide an in-depth understanding of SPEAR, including different pre-training targets (see Appendix G.1), quantiser configuration (see Appendix G.3), masked-prediction pre-training loss (see Appendix G.2), and dual-domain pre-training strategy (see Appendix G.5).
@@ 5.1 Downstream Fine-tuning
We evaluate the performance of the pre-trained models on two key downstream fine-tuning tasks: automatic speech recognition (ASR) for speech capabilities and audio tagging (AT) for general audio understanding capabilities. The downstream fine-tuning results for the single-domain and dual-domain models are presented in Table 4 and configurations are shown in Appendix C.3.
@@
ASR performance is evaluated on LibriSpeech (Panayotov et al., 2015), where the model is fine-tuned on the train-clean-100 subset (LS-100) or the full 960 hours (LS-960) of LibriSpeech. A lightweight, stateless RNN-T decoder (Graves, 2012; Ghodsi et al., 2020) with fewer than 3M parameters using an output vocabulary of 500-class byte-pair-encoding (Sennrich et al., 2016) units is attached to the pre-trained models unless otherwise noted. During fine-tuning, the pre-trained SSL model is also updated. The projection heads for predicting MVQ tokens are discarded. Performance is measured by the Word Error Rate (WER) on the test-clean and test-other splits of LibriSpeech, using beam search decoding with no external language model. We mainly compare with the WavLM models pre-trained on similar corpora at similar model sizes. Additional results of ASR fine-tuning with a CTC (Graves et al., 2006) decoder can be found in Appendix D.
@@
To evaluate audio capabilities, our models are fine-tuned on AudioSet for AT following the procedure in Gong et al. (2021). We perform fine-tuning on both the balanced subset (AS-20k) and the full dataset (AS-2M). A linear projection layer is added on top of the encoder to predict the class probability of 527 sound event classes. Binary cross-entropy is used as the training objective, and the mean average precision (mAP) is measured on the AudioSet evaluation set. We compare SPEAR models with existing state-of-the-art audio SSL models.
| Model | # Params |
LS-100 | LS-960 | AS-20k | AS-2M | ||||
| clean | other | clean | other | |||||||
| Speech SSL Models | ||||||||||
| WavLM Base + △ (Chen et al., 2022) | 95M | 94k | 4.0 | 8.4 | 2.9 | 5.4 | - | - | ||
| HuBERT Large△ (Hsu et al., 2021) | 317M | 60k | - | - | 1.8 | 3.9 | - | - | ||
| WavLM Large△ (Chen et al., 2022) | 317M | 94k | 3.0 | 6.1 | 1.8 | 3.8 | - | - | ||
| Ours, SPEAR Base | 94M | 84k | 3.0 | 5.8 | 1.9 | 4.0 | 26.9 | 43.6 | ||
| Ours, SPEAR Large | 327M | 84k | 2.6 | 4.7 | 1.7 | 3.3 | 26.4 | 43.9 | ||
| Audio SSL Models | ||||||||||
| BEATs (Chen et al., 2023) | 90M | 5k | - | - | - | - | 38.9 | 48.6 | ||
| EAT (Chen et al., 2024) | 88M | 5k | - | - | - | - | 40.2 | 48.6 | ||
| ATST Frame (Li et al., 2024a) | 86M | 5k | - | - | - | - | 39.0 | 48.0 | ||
| Dasheng-Base△ (Dinkel et al., 2024) | 86M | 272k | - | - | - | - | - | 49.7 | ||
| Ours, SPEAR Base | 94M | 13k | 11.2 | 23.0 | - | - | 39.2 | 49.3 | ||
| Ours, SPEAR Large | 327M | 13k | 7.4 | 18.6 | - | - | 39.3 | 49.8 | ||
| Speech & Audio SSL Models | ||||||||||
| Ours, SPEAR Base | 94M | 97k | 3.1 | 6.1 | 1.9 | 4.2 | 39.1 | 48.4 | ||
| Ours, SPEAR Large | 327M | 97k | 2.6 | 4.8 | 1.7 | 3.4 | 39.2 | 49.6 | ||
| Ours, SPEAR XLarge | 600M | 197k | 2.5 | 4.6 | 1.6 | 2.9 | 39.4 | 50.0 | ||
@@ Results
The results presented in Table 4 demonstrate that SPEAR learns high-quality speech and audio representations that transfer well to the downstream tasks. For speech, our SPEAR Base and Large achieve relative WER reductions of 25.9% and 13.2% on the test-other set in LS-960 compared to their WavLM counterparts, with similar model size and pre-training data. For general audio, the SPEAR Base model achieves an mAP of 49.3 on AS-2M, surpassing all other audio SSL models333A strict comparable AT fine-tuning setup with all models pre-trained with the same amount of data is shown in Appendix G.1, where our Base-scaled SPEAR model consistently outperforms SOTA audio SSL models. (except Dasheng Base, which is pre-trained with 20 times more general audio data), and SPEAR Large improves further to 49.8 on AS-2M. These results highlight the effectiveness of using fine-grained MVQ tokens as pre-training targets in SPEAR.
@@
Moreover, the dual-domain models successfully learn a unified representation space capable of handling both tasks with minimal performance loss, achieving ASR and AT performance comparable to their single-domain counterparts, and this performance gap diminishes as model capacity increases. Specifically, the WER for SPEAR Large model on test-other is only 0.1 higher than the speech-domain specialist SPEAR Large, while its mAP is only 0.2 lower than the SPEAR Large. This demonstrates that with sufficient model capacity, our dual-domain pre-training scheme enables a single model to learn a unified representation space with strong capability for both domains. It should be noted that this versatility is particularly important, since the single-domain models yield poor cross-domain capability (see SPEAR on AT or SPEAR on ASR). Finally, our largest model SPEAR XLarge further improves the performance on ASR and AT, setting a new state-of-the-art for SSL models on AS-2M AT task by achieving an mAP of 50.0.
In summary, the experiments in Table 4 suggest that the representations learnt through SPEAR adapt well to both domains after fine-tuning, proving its strong capability of learning both domain-specific and unified representations that excel in both speech and general audio domains.
@@ 5.2 SUPERB Evaluation
Experiments are carried out on SUPERB (Yang et al., 2021; Tsai et al., 2022), a benchmark for evaluating SSL models on a wide range of speech processing tasks. We follow the standard SUPERB evaluation protocol, using a weighted sum of the frozen intermediate representations from the SSL models. For better readability, we group the SUPERB tasks into three categories: Understanding, Paralinguistic, and Enhancement. We select representative tasks within each category and report their results in Table 5. The primary comparison is made against the WavLM Large, which is the current SOTA on SUPERB. Further details on the SUPERB evaluation, including the individual task information and complete results on SUPERB, can be found in Appendix E.
@@ Results
As can be seen from Table 5, SPEAR improves across the range of speech tasks, achieving notable gains across all three task categories. With the same model size and pre-training corpora, our speech-domain SPEAR Large model outperforms the current state-of-the-art WavLM Large on nearly every task. The improvement in paralinguistic capabilities is particularly noteworthy. For instance, SPEAR Large achieves a 16.7% relative reduction of equal-error-rate on speaker verification (SV) and a 1.48% absolute accuracy improvement on emotion recognition (ER). This suggests that our fine-grained masked-token prediction objective helps the model learn richer paralinguistic information beyond speech content alone, than by using k-means clustered tokens. An analysis of the feature subspaces learned through the fine-grained MVQ tokens is conducted in Appendix G.4.
Moreover, our dual-domain SPEAR models maintain strong performance while being able to handle both speech and audio. Despite a slight degradation in some understanding and paralinguistic tasks, the SPEAR Large notably outperforms its speech-only counterpart on keyword spotting (KS) and speech enhancement (SE), indicating a positive synergy from the dual-domain pre-training on these tasks. SPEAR Large also outperforms USAD Large (Chang et al., 2025) comprehensively, another unified speech and audio model trained via matching the representations of two teacher models with more data, demonstrating the advantage of the SSL objective defined by SPEAR. Finally, by scaling up the model size and training data, SPEAR XLarge, our largest dual-domain model, pushes the performance boundary even further, establishing new state-of-the-art on multiple SUPERB tasks, while being more versatile than speech-only models.
| Model | # Param |
Understanding | Paralinguistic | Enhancement | ||||||||||
| PR | ASR | IC | KS | SF | ST | SID | SV | ER | SE | SS | |||||
| Speech SSL models | |||||||||||||||
| WavLM Base+ | 95M | 94k | 3.5 | 3.92 | 99.00 | 97.37 | 90.6 | 24.3 | 89.4 | 4.07 | 68.7 | 2.63 | 10.85 | ||
| WavLM Large | 317M | 94k | 3.1 | 3.44 | 99.31 | 97.86 | 92.2 | 26.6 | 95.5 | 3.77 | 70.6 | 2.70 | 11.19 | ||
| Ours, SPEAR Base | 94M | 84k | 3.4 | 3.46 | 99.17 | 97.50 | 91.0 | 24.4 | 90.5 | 3.75 | 69.2 | 2.64 | 10.84 | ||
| Ours, SPEAR Large | 327M | 84k | 2.6 | 3.27 | 99.47 | 97.89 | 92.8 | 26.2 | 95.5 | 3.14 | 72.1 | 2.71 | 11.20 | ||
| Audio SSL models | |||||||||||||||
| BEATs | 90M | 5k | 36.4 | 36.4 | 97.70 | 53.40 | - | - | 57.1 | - | 64.5 | - | - | ||
| EAT | 88M | 5k | 55.0 | 25.9 | 92.80 | 62.50 | - | - | 45.0 | - | 62.5 | - | - | ||
| ATST Frame | 86M | 5k | 20.4 | 18.8 | 95.10 | 85.40 | - | - | 69.8 | - | 64.4 | - | - | ||
| Speech+Audio Models | |||||||||||||||
| USAD Base | 94M | 126k | 5.1 | 7.70 | 98.30 | 97.10 | - | - | 88.6 | - | 68.0 | - | - | ||
| USAD Large | 330M | 126k | 4.0 | 6.50 | 98.40 | 97.10 | - | - | 91.2 | - | 68.4 | - | - | ||
| Ours, SPEAR Base | 94M | 97k | 3.9 | 3.76 | 98.05 | 97.58 | 90.5 | 24.1 | 90.0 | 3.85 | 69.4 | 2.66 | 10.89 | ||
| Ours, SPEAR Large | 327M | 97k | 3.1 | 3.39 | 99.40 | 97.92 | 92.1 | 25.6 | 95.0 | 3.30 | 71.6 | 2.72 | 11.12 | ||
| Ours, SPEAR XLarge | 600M | 97k | 2.9 | 3.19 | 99.61 | 98.12 | 92.9 | 26.7 | 96.3 | 2.86 | 73.3 | 2.72 | 11.24 | ||

### v2

@@ 4 Experimental Setup
| Domain | Dataset | Hours |
| Speech | Libriheavy (Kang et al., 2024) | 50k |
| GigaSpeech (Chen et al., 2021) | 10k | |
| VoxPopuli (en) (Wang et al., 2021) | 24k | |
| Yodas-granary (en) (Koluguri et al., 2025) | 100k | |
| Audio | AudioSet (Gemmeke et al., 2017) | 5k |
| VGGsound (Chen et al., 2020) | 0.5k | |
| Freesound (Wu et al., 2023) | 2.8k | |
| Music4all (Santana et al., 2020) | 1k | |
| MTG-Jamendo (Bogdanov et al., 2019) | 3.8k |
| Model | Domain(s) | Data Groups | Total Hours |
@@
| SPEAR-{Base, Large} | Speech & Audio | SA-97k | 97k |
| SPEAR XLarge | Speech & Audio | SA-197k | 197k |
Dataset
Pre-training is performed on both public unlabelled English speech datasets and general audio datasets, as shown in Table 1. Due to the limited amount of publicly available general audio datasets, we incorporate two music datasets, Music4all and MTG-Jamendo, to enrich the general audio data. The pre-training data used for different scales of SPEAR models are given in Table 2.
Model Architecture
As shown in (Shi et al., 2024a), modelling speech representations at different time resolutions improves speech SSL models. Therefore, Zipformer (Yao et al., 2024) is selected as the feature encoder in SPEAR due to its dynamic downsampling mechanism in the intermedi- ate layers. The model receives 128-dimensional filter-bank features as input and produces frame-level representations at a 50 Hz frame rate. We pre-train three model sizes: Base (94M), Large (327M), and XLarge (600M). Detail regarding the encoder configurations is provided in Appendix B.1.
@@
At the Base and Large scales, we pre-train both single-domain and dual-domain models. For the XLarge scale, we only train a dual-domain model with all the data. The pre-training data used for different scales of SPEAR models are given in Table 2. The training data groups are defined as follows: Speech-84k comprises Libriheavy, GigaSpeech, and VoxPopuli (en); Audio-13k includes all general audio datasets in Table 1; SA-97k combines Speech-84k and Audio-13k; and SA-197k additionally includes Yodas-granary. Token mixing is applied to 10% of all training samples, where the mixing signal is sampled from the same batch. Before mixing the waveforms, we sample an SNR between (-5, 5) and scale the second signal accordingly to match the SNR. Both signals are mixed across their full duration with a random offset. Other hyperparameters for pre-training are given in Appendix B.2.
Teacher Models and MVQ
WavLM Large (Chen et al., 2022) and Dasheng 1.2B (Dinkel et al., 2024) are used to generate pre-training targets for speech and audio domains in the main paper, respectively. WavLM Large is a speech SSL model pre-trained on 94k hours of unlabelled English speech data, including Libri-light (Kahn et al., 2020), GigaSpeech, and Voxpopuli. It can be fairly compared with SPEAR models trained on Speech-84k since Libriheavy is the segmented version of Libri-light. The model generates a 1024-d feature at a 50 Hz frame rate. The speech MVQ quantiser with 16 codebooks is trained using the 21st layer representations of WavLM Large. Dasheng 1.2B is pre-trained with an MAE objective on an exceptionally large amount of speech and audio data of over 272k hours, roughly half of which is non-speech general audio. This model generates 1536-d features at 25 Hz. The audio MVQ quantiser is trained on the Dasheng final-layer representations with 8 codebooks. Ablation studies on different teacher models for pre-training target generation and quantiser configurations are presented in Appendix F.1.1 and Appendix F.5.
5 Experimental Results
To validate the effectiveness of the SPEAR framework, we assess its performance through both full fine-tuning and frozen representation evaluations on two major benchmarks for evaluating speech and audio representations: SUPERB (Yang et al., 2021; Tsai et al., 2022) and HEAR (Turian et al., 2022). Finally, ablation studies on core components of SPEAR are discussed in Section 5.4.
@@ 5.1 Downstream Fine-tuning
We evaluate the performance of SPEAR on two key downstream fine-tuning tasks: automatic speech recognition (ASR) for speech capabilities and audio tagging (AT) for general audio understanding capabilities. The results are presented in Table 3 and detailed fine-tuning configurations are shown in Appendix B.3.
@@
The ASR performance is evaluated on LibriSpeech (Panayotov et al., 2015), where the model is fine-tuned on the train-clean-100 subset (LS-100) or the full 960 hours (LS-960). The ASR decoder is a stateless RNN-T decoder (Ghodsi et al., 2020) with 3M parameters and a 500-unit BPE vocabulary (Sennrich et al., 2016). Performance is measured by the Word Error Rate (WER) on the test-clean and test-other sets of LibriSpeech, using beam search decoding with no external language model. Additional results of fine-tuning with a CTC (Graves et al., 2006) decoder are shown in Appendix C.
@@
To evaluate audio capabilities, our models are fine-tuned on AudioSet for AT following the procedure in Gong et al. (2021). We perform fine-tuning on the balanced subset (AS-20k) and the full dataset (AS-2M). A linear projection layer is added after the encoder to predict the probability of 527 sound event classes. Binary cross-entropy is used for optimisation, and the mean average precision (mAP) is measured on the AudioSet evaluation set.
@@ Results
| Model | # Params | Pre-train data | LS-100 | LS-960 | AS-20k | AS-2M | ||
| clean | other | clean | other | |||||
| Speech SSL Models | ||||||||
| WavLM Base + △ | 95M | 94k | 4.0 | 8.4 | 2.9 | 5.4 | - | - |
| WavLM Large△ | 317M | 94k | 3.0 | 6.1 | 1.8 | 3.8 | - | - |
| Ours, SPEAR Base | 94M | 84k | 3.0 | 5.7 | 1.9 | 4.0 | 26.9 | 43.6 |
| Ours, SPEAR Large | 327M | 84k | 2.6 | 4.7 | 1.7 | 3.3 | 26.4 | 43.9 |
| Audio SSL Models (AudioSet 5k hours) | ||||||||
| BEATs (Chen et al., 2023) | 90M | 5k | - | - | - | - | 38.9 | 48.6 |
| EAT (Chen et al., 2024a) | 88M | 5k | - | - | - | - | 40.2 | 48.6 |
| ATST Frame (Li et al., 2024) | 86M | 5k | - | - | - | - | 39.0 | 48.0 |
| Ours, SPEAR Base | 94M | 5k | - | - | - | - | 39.0 | 49.3 |
| Ours, SPEAR Large | 327M | 5k | - | - | - | - | 39.3 | 49.7 |
| Ours, SPEAR Base | 94M | 13k | 11.2 | 23.0 | - | - | 39.2 | 49.3 |
| Ours, SPEAR Large | 327M | 13k | 7.4 | 18.6 | - | - | 39.3 | 49.8 |
| Speech & Audio SSL Models | ||||||||
| USAD Base△ | 94M | 126k | 4.9 | 11.4 | - | - | 35.7 | - |
| USAD Large△ | 330M | 126k | 4.0 | 7.7 | - | - | 37.4 | - |
| Ours, SPEAR Base | 94M | 97k | 3.1 | 6.0 | 1.9 | 4.2 | 39.1 | 48.6 |
| Ours, SPEAR Large | 327M | 97k | 2.6 | 4.8 | 1.7 | 3.4 | 39.2 | 49.6 |
| Ours, SPEAR XLarge | 600M | 197k | 2.4 | 4.6 | 1.6 | 2.9 | 39.4 | 50.0 |
| Model | # Param | Pre-train data | Understanding | Paralinguistic | Enhancement | |||||||
| PR | ASR | IC | KS | SID | SV | SD | ER | SE | SS | |||
| Speech SSL models | ||||||||||||
| WavLM Base+ | 95M | 94k | 3.5 | 3.92 | 99.00 | 97.37 | 89.4 | 4.07 | 3.5 | 68.7 | 2.63 | 10.85 |
| WavLM Large | 317M | 94k | 3.1 | 3.44 | 99.31 | 97.86 | 95.5 | 3.77 | 3.2 | 70.6 | 2.70 | 11.19 |
| Ours, SPEAR Base | 94M | 84k | 3.4 | 3.43 | 99.17 | 97.50 | 90.5 | 3.77 | 2.9 | 69.1 | 2.64 | 11.32 |
| Ours, SPEAR Large | 327M | 84k | 2.6 | 3.24 | 99.44 | 97.88 | 95.6 | 3.14 | 2.3 | 72.1 | 2.72 | 11.67 |
| Audio SSL models | ||||||||||||
| BEATs | 90M | 5k | 36.4 | 36.4 | 97.70 | 53.40 | 57.1 | - | 5.2 | 64.5 | - | - |
| EAT | 88M | 5k | 55.0 | 25.9 | 92.80 | 62.50 | 45.0 | - | 4.7 | 62.5 | - | - |
| Dasheng 1.2B | 1.2B | 272k | 14.3 | 13.8 | 98.13 | 97.73 | 92.4 | - | 3.8 | 68.7 | - | - |
| Speech & Audio Models | ||||||||||||
| USAD Base | 94M | 126k | 5.1 | 7.70 | 98.30 | 97.10 | 88.6 | - | 4.2 | 68.0 | - | - |
| USAD Large | 330M | 126k | 4.0 | 6.50 | 98.40 | 97.10 | 91.2 | - | 3.9 | 68.4 | - | - |
| Ours, SPEAR Base | 94M | 97k | 3.8 | 3.79 | 98.13 | 97.59 | 90.3 | 3.85 | 3.0 | 69.2 | 2.67 | 11.39 |
| Ours, SPEAR Large | 327M | 97k | 3.1 | 3.39 | 99.40 | 97.92 | 95.2 | 3.30 | 2.3 | 71.7 | 2.73 | 11.60 |
| Ours, SPEAR XLarge | 600M | 197k | 2.9 | 3.15 | 99.60 | 98.12 | 96.3 | 2.86 | 2.0 | 73.3 | 2.75 | 11.73 |
@@
The results presented in Table 3 demonstrate that SPEAR learns high-quality representations in both single-domain and dual-domain setups. In the speech domain, SPEAR Base and Large achieve relative WER reductions of 25.9% and 13.2% on the test-other set in LS-960 compared to their WavLM counterparts with similar model size and pre-training data. In the audio domain, SPEAR Base pre-trained on AudioSet achieves an mAP of 49.3 on AS-2M, surpassing all other audio SSL models pre-trained on AudioSet. After increasing the model size, SPEAR Large improves further to 49.7 on AS-2M, setting a new state-of-the-art for SSL models pre-trained solely on AudioSet.
@@
Moreover, our dual-domain models successfully learn a unified representation space capable of handling both tasks with minimal performance loss comparable to their single-domain counterparts, and this performance gap diminishes as model capacity increases. Specifically, the WER for SPEAR Large model on test-other is only 0.1 higher than the speech-domain specialist SPEAR Large, while its mAP is only 0.2 lower than the SPEAR Large. It should be noted that this versatility is particularly important, since all single-domain models show poor cross-domain transfer capability (see SPEAR on AT and SPEAR on ASR). Compared to USAD, SPEAR consistently achieves better performance while using less pre-training data.
@@ 5.2 SUPERB Evaluation
Experiments are carried out on SUPERB (Yang et al., 2021; Tsai et al., 2022), a benchmark for evaluating SSL representations on a wide range of speech processing tasks. We follow the standard SUPERB evaluation protocol, using a weighted sum of the frozen intermediate representations from the SSL models. For better readability, we group the SUPERB tasks into three categories: Understanding, Paralinguistic, and Enhancement. We select representative tasks within each category and report their results in Table 4. The primary comparison is made against WavLM Large, which is the current SOTA on SUPERB. We also compare our dual-domain models SPEAR with USAD, another unified speech and audio model trained via multi-teacher feature matching. Further details and complete results on SUPERB can be found in Appendix D.
@@ Results
As can be seen from Table 4, SPEAR improves across the range of speech tasks, achieving notable gains across all three task categories compared to the current SOTA WavLM. Specifically, with the same model size and pre-training corpora, the speech domain SPEAR Large model outperforms the current SOTA WavLM Large on nearly every task. The improvement in paralinguistic capabilities is particularly noteworthy: SPEAR Large achieves a 16.7% relative reduction of equal-error-rate on speaker verification (SV) and a 1.5% absolute accuracy improvement on emotion recognition (ER). This suggests that masked prediction of fine-grained MVQ tokens in SPEAR helps the model learn richer paralinguistic information beyond lexical content alone than by using k-means tokens. An analysis of the feature subspaces learned through the MVQ tokens is conducted in Appendix F.4.2.
Our dual-domain SPEAR models maintain strong performance on SUPERB while being more versatile than speech-domain models. Despite a slight degradation in some understanding and paralinguistic tasks, SPEAR Large outperforms its speech-only counterpart SPEAR Large on keyword spotting (KS) and speech enhancement (SE), indicating a positive synergy from the dual-domain pre-training. SPEAR also consistently outperforms USAD at comparable model sizes while using less training data. After scaling, SPEAR XLarge pushes the performance boundary even further, establishing new SOTA on most tasks.
It is also evident that audio SSL models generally underperform on the SUPERB benchmark. Even with a very large pre-training corpus of 272k hours containing both speech and general audio data, Dasheng 1.2B consistently performs more poorly than our much smaller SPEAR or SPEAR Large, especially on understanding-based tasks (e.g., ASR). This gap could be attributed to the MAE objective used by Dasheng, which tends to focus on low-level acoustic detail rather than high-level semantic structures necessary for speech understanding. In contrast, by leveraging fine-grained MVQ tokens extracted from domain-specific teachers, SPEAR allows the model to learn semantic structures while retaining acoustic details through masked-token prediction, leading to a comprehensive performance over both domains after pre-training on speech and audio data.